Use the properties of inverse functions to evaluate the expression
cos [arcsin (5sqrt29/29)]
I was told the answer is 2sqrt29/29 but I can't figure out how. Thanks!

Answer by stanbon(48535)

evaluate the expression
cos [arcsin (5sqrt29/29)]
---------------------------------
Sketch a rt. triangle in standard position in the 1st Quadrant.
You have an angle whose sine is (5sqrt(29))/29
So the hypotenuse is 29 and the vertical side is 5sqrt(29)
---------------
Use Pythagoras to solve for the horizontal side (hs)
(hs)^2 = 29^2 -(5sqrt(29))^2
(hs)^2 = 841 - 725
(hs) = 10.77...
--------------------
Therefore cos[angle] = 10.77/29
===================================
Cheers,
Stan H.